On average per month, a particular laboratory uses 15 points to identify and match a fingerprint from a crime scene with a suspect in the database. Calculate the following probabilities:
a. The probability of a successful match in a given month with either 12 or 17 points of identification.
1
Expert's answer
2020-06-16T18:51:59-0400

"P(Z) = \frac {x- \mu} {\sigma}"


"= p(\\frac {12- 15} {1})+p(\\frac {17- 15} {1})=p(-3)+p(2) = 0.00135+0.97725 =0.9786\n\u200b"